A solid shaft 100 mm in diameter is subjected simultaneously to an axial compressive force of 50 kN and to a torque that twist through an angle of 1.5
1
Expert's answer
2021-07-12T03:36:39-0400

If we were to find the torque

If "G_{steel} = 85 GPa" we can find the torque of the twist.

"\frac{T}{J}=\frac{G \theta}{L} \implies T= \frac{G \theta}{L}J\\
T= \frac{ 85*10^9*\frac{1.5\pi}{180}}{1}*\frac{\pi}{32}*0.1^4\\
T=21.84678 kNm"